OpenStudy (anonymous):

Find the limit of sinx/x^3 if x approaches to 0-.

OpenStudy (dape):

This limit diverges to positive infinity. In this case, as x->0, sin(x) can be replaced by x. So we get \[ \lim_{x->0^-}=\lim_{x->0^-}\frac{\sin x}{x^3}=\lim_{x->0^-}\frac{1}{x^2}=+\infty\]

OpenStudy (anonymous):

How did you get 1/x^2?

OpenStudy (dape):

As sin(x) approaches 0, the approximation sin(x)=x becomes better and better, so I simply got 1/x^2 by replacing sin(x) by x.

zepdrix (zepdrix):

If that's too confusing Ideal, consider that this problem is a good candidate for L'Hopital's Rule :o

OpenStudy (anonymous):

I did L'Hopital's Rule and got -1/6, but that's not the answer.

OpenStudy (dape):

Indeed, in general when you have 0/0 as the "limit" you can use L'Hopital's Rule: \[ \lim_{x\rightarrow0}\frac{\sin x}{x^3}=\lim_{x\rightarrow0}\frac{\cos x}{3x^2} \] The numerator tends to 1, a finite value, the denominator tends to 0, which means that the limit blows up, i.e. diverges to infinity.

zepdrix (zepdrix):

\[\Large L'H \to\qquad=\lim_{x\to0^-}\frac{\cos x}{3x^2}\]After you apply L'Hop once, you shouldn't be able to perform it again. We're no longer getting an indeterminate form that allows us to perform it again.

OpenStudy (anonymous):

It's not related to ode, my prof gave us this assignment to refresh our brains.

OpenStudy (anonymous):

Thanks guys.

OpenStudy (anonymous):

this would also be an excellent candidate for a power series representation you would get the answer instantly

OpenStudy (loser66):

@Idealist here. I need place to answer, right?

OpenStudy (loser66):

do as they teach you: L'hopital rule. do you know that rule? just take derivative both numerator and denominator. if the result still form the form of 0/0 or infty/infty , take l'hopital rule again. ok?

OpenStudy (anonymous):

Okay.

OpenStudy (loser66):

show me your work, stubborn girl. it's not hard.

OpenStudy (anonymous):

It's in my new post.

OpenStudy (loser66):

where? just tag me

OpenStudy (loser66):

oh, I saw it, but too many people there. I don't show up